OpenStudy (anonymous):

can someone solve this for me? ive been on this question fr days.. 5x+10/x*2 * 2x/4x-10 multipy and solve

OpenStudy (jhannybean):

Have you tried solving it? I've showed you how. Revisit your other post and tell em what you do not understand.

OpenStudy (anonymous):

I've tried to. I don't understand how to solve it.

OpenStudy (babyx3boo):

is there any multiple choice?

OpenStudy (anonymous):

no I jus have to solve it

OpenStudy (anonymous):

I've got it down to 5•2x/4x-10 but I can't solve the rest

OpenStudy (babyx3boo):

is it 2x or x squared?

OpenStudy (anonymous):

it's 5x+10/x+2 • 2x/4x-10

OpenStudy (jhannybean):

\[\large \frac{5x+10}{x+2} \cdot \frac{2x}{4x-10}\]How can you simplify \(5x+10\)?

OpenStudy (babyx3boo):

Does this help?

OpenStudy (anonymous):

I honestly don't know!

OpenStudy (anonymous):

I have it down to 5•2x/4x-10

OpenStudy (babyx3boo):

check the picture that i send you.

OpenStudy (anonymous):

idk I'm so damn confused by this.. I've asked for so much help and I still don't understand it

OpenStudy (jhannybean):

Well, you're on the right path. \[\frac{5(2x)}{4x-10}\] Now just multiply \(5 \cdot 2x\) . Wht does that give you.

OpenStudy (anonymous):

10x?

OpenStudy (jhannybean):

yep. so you've now got \[\frac{10x}{4x-10}\] Now so you see that \(4x-10\) has a common LCM of 2, since 2(2) = 4, and 2(5) =10.

OpenStudy (jhannybean):

We pull out a common multiple of 2. \[4x-10 \iff 2(2x+5)\]

OpenStudy (anonymous):

okay I understand that

OpenStudy (anonymous):

I don't understand the second part

OpenStudy (jhannybean):

So, like @babyx3boo illustrated, you will have \[\frac{\color{red}{10}x}{\color{red}{2}(2x+5)}\]What is \(\frac{10}{2}\)?

OpenStudy (anonymous):

would the answer be 2??

OpenStudy (anonymous):

10/2 would be 5?

OpenStudy (anonymous):

@jhannybean

OpenStudy (anonymous):

me too help me